# Reformulating Sequential Recommendation: Learning Dynamic User Interests with Content-enriched Language Modeling

### v2

@@ Abstract
Recommender systems are essential for online applications, and sequential recommendation has enjoyed significant prevalence due to its expressive ability to capture dynamic user interests. However, previous sequential modeling methods still have limitations in capturing contextual information. The primary reason for this issue is that language models often lack an understanding of domain-specific knowledge and item-related textual content. To address this issue, we adopt a new sequential recommendation paradigm and propose LANCER, which leverages the semantic understanding capabilities of pre-trained language models to generate personalized recommendations. Our approach bridges the gap between language models and recommender systems, resulting in more human-like recommendations. We demonstrate the effectiveness of our approach through experiments on several benchmark datasets, showing promising results and providing valuable insights into the influence of our model on sequential recommendation tasks. Furthermore, our experimental codes are publicly available.
@@ 1 Introduction
Recommender systems are crucial engines of various online applications, including e-commerce, advertising, and online videos. They play a pivotal role in discovering user interests and alleviating information overload, thus assisting users in their decision-making process. Over the past years, collaborative filtering (CF) [20] has become one of the most prevalent techniques due to its effectiveness and efficiency in large-scale recommendation scenarios.
As one type of model-based CF technique, the matrix factorization algorithm [11] has been recognized as one of the most successful techniques in the recommendation community. Despite its effectiveness, this approach is not good at capturing dynamic user interests since temporal properties are largely overlooked in user-item interaction matrix. Later, a series of works are devoted to modeling evolving user preference. Among them, sequential recommendation [19, 7, 1] has become increasingly prevalent since the sequence of historical behavior ordered by interacted time can clearly reflect each user’s long-term and short-term preference, as shown in Figure 1(a).
@@
By carefully revisiting recent progress in the sequential recommendation, we notice much of the great progress has been driven by achievements in network architecture or pre-training algorithms in natural language processing [5]. For one thing, a large body of advanced architecture models [9, 17, 27] has facilitated the building of high-capacity recommendation models. For another, some researchers attempt to leverage the powerful capacity of language modeling in semantic understanding to perform content-based sequential recommendations [29, 26]. In addition, we also notice a few related advancements attempt to integrate pre-trained language models (PLMs) into recommender systems, leveraging their ability to capture semantic information and achieving promising results [14]. However, these works do not depart from the previous paradigm, merely adding a text encoder and continuing to use traditional sequential modeling methods, as depicted in Figure 1(b).
@@
With the continuous improvement of model capabilities, we ponder on whether the paradigm of sequential recommendation can be revolutionized to recommend items to users in a way that mimics human reasoning. Some recent works have attempted to use powerful PLMs as recommender systems by formulating recommendation as a language modeling task, with the item IDs [6] or titles [31] as input to the language model. Despite leveraging the inferential capabilities of language models for generating sequential recommendations, these methods have yielded unsatisfactory performance, falling behind the previous state-of-the-art results.
@@
In light of the above, language models perform suboptimally to sequential recommendation systems is their lack of domain-specific knowledge and mastery of item-related textual content. For example, in movie recommendation, The Matrix could be understood by the language model as a mathematical matrix, rather than being associated with hacker-related information, which would significantly reduce the effectiveness of the recommendation. Moreover, language models are limited by the scarcity of pre-training data and the constraint of input text length during inference, which makes it challenging to incorporate sufficient item content knowledge to facilitate recommendations.
@@
To address these issues, we propose LANguage modeling for Content-Enriched Recommendation (LANCER) to employ PLMs as sequential recommender by integrating domain-specific knowledge and user behavior. Specifically, we introduce a knowledge prompt that facilitates the acquisition of domain-specific knowledge, and a reasoning prompt that integrates the acquired knowledge and item content information to generate personalized recommendations. By transforming the sequential recommendation into a text generation task, we change the paradigm of sequential recommendation, and Figure 1(c) illustrates the new paradigm.
Figure 1: A paradigm shift in sequential recommender systems.
@@
  • We propose LANCER that incorporates domain knowledge and item content prompts into PLMs, leveraging their semantic understanding abilities to generate user recommendation results. This is a new paradigm for sequential recommendation systems.
  • Our designed model offers the advantage of integrating external knowledge to enhance the modeling of item content understanding. Our approach bridges the gap between language models and recommender systems by fusing knowledge from both domains.
@@
2 Preliminaries
Consider a sequential recommendation task, where the input is the items interacted by a user, formulated by a sequence of items $i_{1},i_{2},...,i_{n-1}$, and the output is the prediction of the next item $i_{n}$ that the user may interest. Traditional methods encode each item as a vector and use sequential modeling methods to predict the next item using classification layer, and calculate the result through the predicted probability distribution $p(i_{n}\mid i_{1},i_{2},...,i_{n-1})$.
When it comes to language models as sequential recommender system, we need to change the above definition. In our work, we represent each item in a natural language way rather than using a discrete representation (such as ID). Specifically, we utilize information consistent with human perception, such as using the movie title. For item $i_{k}$, we map it to encoded tokens $e_{k}=[t_{k_{1}},t_{k_{2}},...,t_{k_{m}}]$. We use a prompt $f(\cdot)$ to convert the historical sequence of interactions between users and items into natural language text $s=f([e_{1},e_{2},\ldots,e_{n-1}])$ as context input to the model. Then we need the model to generate several tokens and decode these tokens into natural language, representing the next possible items the user may interact with. The output of the model will be in natural language, thus the probability distribution we need to calculate is equation 1 until the model generates the $[EOS]$ (End Of Sequence) token.
$p(e_{n}\mid s)=\prod_{j}p(t_{n_{j}}\mid t_{n_{<j}},s).$ (1)
We adopt a transformer-based autoregressive PLM to validate our method. The interaction between users and items is unidirectional sequential data arranged in chronological order, making the autoregressive model suitable for modeling this type of problem. We utilize the autoregressive nature of the model to predict the next token based on preceding tokens. This prediction task is achieved by minimizing the negative log-likelihood of the subsequent token:
$\mathcal{L}=\sum_{n}-\log p\left(t_{n}\mid t_{1},\ldots,t_{n-1};\Theta_{LM}\right),$ (2)
where $t_{n}$ is the next token to be predicted, and $\Theta_{LM}$ represents all parameters in Transformer.
@@ 3 Methodology
In this section, we introduce the structure of LANCER, elaborating on its input/output and basic principles. In Section 3.1, we introduce the framework and operation of the model. In Section 3.2, we describe how we use a knowledge prompt to learn domain-specific knowledge in the recommendation domain and obtain vector representations for each item. In Section 3.3, we explain how our designed reasoning prompt integrates a user’s interests, historical behavior and domain-specific knowledge to enable language model-based recommendations. Finally, in Section 3.4, we elaborate on how to map the text generated by our model to a specific item for recommendation.
@@ 3.1 Model Architecture
The proposed LANCER architecture is shown in Figure 2. For each item in the user interaction history, we extract all text content information and input it into a language model with fixed parameters to train the knowledge prompt and obtain a vector representation $h_{i}$ for each item. Afterward, we use an attention mechanism to combine domain knowledge with history of user interaction as the reasoning prompt, allowing the model to generate the next item that the user may be interested in. Finally, the generated text is mapped to one of the candidate items recommended to the user.
3.2 Knowledge Prompt
Once directly applying a language model for item recommendation, a major challenge arises as language models may not possess the knowledge of the relevant domains. Previous works [31, 6, 3] attempt to use movie IDs or movie titles as input, relying on language model to make recommendations. However, these approaches fail to fully harness the text understanding capabilities of PLMs. Movie IDs, being numerical information, can not convey the underlying meaning to the language model, and even movie titles may not accurately reflect the genre and details of the movie.
@@ 3.2 Knowledge Prompt
To address the above issues, we design a continuous prompt to learn recommendation domain information while avoiding catastrophic forgetting that may occur with language model retraining. Inspired by Prefix-tuning [13] and P-tuning [16, 15], we add continuous prefix prompts to guide the PLM without changing the parameters. We define this as a knowledge prompt. The knowledge prompt adds parameters in front of each layer of the language model, and the effect of these continuous parameters propagates to the whole transformer activation layer and then to subsequent tokens. In terms of implementation, we set knowledge tokens with a fixed length and corresponding embedding spaces, and then generate the hidden states of each layer’s attention blocks in the transformer through a fully connected layer. We do not directly optimize the related parameters because updating the prompt parameters directly leads to unstable optimization and a slight drop in performance [13]. Our calculation process is as follows:
$KP=MLP(Emb([k_{1},\ldots,k_{\theta}])),$ (3)
where $k_{i}$ denotes $i$-th knowledge token, $\theta$ denotes the number of knowledge tokens. Suppose a PLM with $L$ layers, for an attention computation in the $l$-th layer, a distinct prefix of key-value pairs $K_{l}$ and $V_{l}$ is learnt, which are augmented to become:
$K_{l}^{{}^{\prime}}=[KP_{l,K};K_{l}],\quad V_{l}^{{}^{\prime}}=[KP_{l,V};V_{l}],$ (4)
@@
where $K_{l}^{{}^{\prime}},V_{l}^{{}^{\prime}}\in\mathbb{R}^{(\theta+M)\times d}$, $M$ denotes the input length of the language model, and $d$ denotes the dimension of hidden layers.
3.3 Reasoning Prompt
With the help of knowledge prompts, we can provide more accurate recommendations for a specific domain. However, this method lacks a modeling of user behavior and a deeper understanding of item information. Using only knowledge prompts, the model obtains the same domain information for different users. Yet, user preferences often lie within a more personalized subset of domain information, rather than the entire domain. For instance, a user who enjoys science fiction movies may benefit from a knowledge prompt that selects more science fiction items based on their viewing history, resulting in more accurate recommendations.
To address this issue, we propose a method called reasoning prompt to learn dynamic user interest. Reasoning prompt combines domain knowledge and user behavior to generate prompts for the language model to provide personalized recommendations. Specifically, reasoning prompt consists of two parts: first, the embeddings of knowledge tokens are mapped to the embedding space and then shared with the knowledge prompt via a fully connected layer to obtain a hidden state; second, for each item, we use the last hidden state $h_{i}$ of the last input token of computed by language model with the knowledge prompt, which considers all previous tokens, to represent the item $i$. Equation 5 shows how the hidden layer is calculated:
$h_{i}=\Theta_{LM}(\mathcal{D}\mid K_{l}^{{}^{\prime}},V_{l}^{{}^{\prime}}),$ (5)
@@ 3.3 Reasoning Prompt
where $\mathcal{D}$ denotes the item $i$ detailed text information. Therefore, a user’s behavior sequence can be expressed as $[h_{1},\ldots,h_{t-1}]$. We will pad if the length of the sequence less than the fixed $N$.
@@
Next, we use an attention mechanism to combine domain knowledge and user behavior. To obtain user-specific domain information, we extract specialized domain information based on the user’s interest history, which can help to provide more effective recommendations. We use the user’s historical behavior representation as query, domain information as key and value, and extract the unique prompt prefix for each user to help the subsequent model to better recommend:
$RP=MLP({softmax}\left(\frac{\mathcal{Q}\mathcal{K}^{T}}{\sqrt{d_{k}}}\right)\mathcal{V}).$ (6)
In equation 6, $\mathcal{Q},\mathcal{K},\mathcal{V}$ denotes query, key and value of attention mechanism respectively. They are calculated as follows:
$\mathcal{Q}=\mathbf{w}_{q}[h_{1},\ldots,h_{t-1}]+\mathbf{b}_{q},$ (7)
$\mathcal{K}=\mathbf{w}_{k}KP+\mathbf{b}_{k},$ (8)
$\mathcal{V}=\mathbf{w}_{v}KP+\mathbf{b}_{v},$ (9)
where $\mathcal{Q}\in\mathbb{R}^{N\times d}$, $\mathcal{K},\mathcal{V}\in\mathbb{R}^{\theta\times d}$, and $\mathbf{w}$ and $\mathbf{b}$ denote the weight and bias parameters of the trainable linear layer separately.
3.4 Item Mapping
@@
In the generation phase, we use beam search decoding to generate output sequences with the 10 highest scores, then truncate each output to include only the first new item. Beam search decoding at step $t$ can be expressed as:
$E_{t}=\operatorname*{arg\,max}_{E\in\mathcal{C}_{t},|E|=w}\log p(E\mid\Theta_{LM}(s,RP)),$ (10)
$\mathcal{C}_{t}=\{[e_{t-1},e]\ |\ e\in\mathcal{V},e_{t-1}\in E_{t-1}\},$ (11)
where $\mathcal{V}$ is the vocabulary set, $w$ is the number of beams, and $E$ is the current set of top-$w$ outputs.
At this point, we have obtained the results of the model prediction, but in order to calculate the accuracy of our prediction, we need to map the predicted result to one of all items in database. Each output is matched to an item in the database according to the following procedure. First, the text sequence is passed through the embedding layer of the autoregressive model, resulting in a sequence of token embeddings. We proceed to perform max pooling over the token embeddings to obtain a fixed length vector representation of the output. Finally, we compute the cosine similarity in equation 12 between the vector representations of the output $i_{pred}$ and the titles of each item in the database $\mathcal{I}$, then identify the item with the highest similarity. The representations of the total set of items can be computed only once and used throughout the generation process.
$i^{{}^{\prime}}=\mathop{\arg\max}_{k\in\mathcal{I}}CosSim(i_{pred},i_{k}).$ (12)
@@ 4 Experiments
In this section, we first describe the experimental setup and compare them with other competitive baselines. We then analyze the experimental results of recommendation performance.
@@ 4.1 Experimental Setup
We conduct experiments on three public datasets:
  • MovieLens [^1] https://grouplens.org/datasets/movielens/ is a widely used movie recommendation dataset. For each user, we order their rated movies chronologically and use the movie name as a movie’s representation. Specifically, we conduct experiments on a 1M-sized dataset.
  • MIND [^2] https://msnews.github.io/ [28] is a dataset for English news recommendation. We use the provided historical interactions to construct user behavior sequences and use the title of the news article to represent each item. In this paper we use the small version of MIND dataset.
  • Goodreads [^3] https://sites.google.com/eng.ucsd.edu/ucsdbookgraph/home [25, 24] datasets contain reviews from the Goodreads book review website, and a variety of attributes describing the items. For each user, we order the books they added to their public shelf according to interaction time to obtain historical sequences, and use the title to denote each book. We use a subset on poetry from Goodreads dataset.
For all three datasets, we limit the sequence length to between 5 and 10 items, with sequences of less than 5 items are omitted and sequences of more than 10 items are truncated from the beginning to include only the last 10 items. Detailed statistics of the processed datasets are presented in Table 1.
@@
In this work, Recall and NDCG are employed as evaluation metrics. The former is an evaluation of un-ranked retrieval sets while the latter reflects the order of ranked lists. Here, we consider top-N (e.g., N = 5, 10) for recommendations. We follow the leave one-out strategy described in [9] to split the interactions into training, validation, and testing sets. For the testing stage, we adopt the all-ranking protocol to evaluate recommendation performance. To evaluate our model, we compute Recall and NDCG for the top 5 and 10 items respectively.
@@
Dataset Num. User Num. Item Num. Inteaction Sparsity Avg. Actions/User Avg. Actions/Item
MovieLens 6040 3231 72480 99.63% 12.0 22.43
MIND 91935 44908 2248027 99.94% 24.45 50.06
Goodreads 120968 28480 1095758 99.97% 9.06 38.47
@@
We use the GPT-2 [18] pre-trained model with 124M parameters from Hugging Face  [^4] https://huggingface.co/gpt2 as our backbone model for experiments. During training, we employ the Adam [10] optimizer with a linear schedule learning rate decay. For different datasets, we keep the following hyperparameters consistent: batch size of 16, maximum token length of 512 for the language model input, maximum input length of 32 tokens for each item (truncated for items exceeding the limit). Due to the varying sizes of the datasets, there are differences in the learning rate and epochs. We employ the Sequential Tuning approach from [12], performing Fixed-LM Prompt Tuning and Prompt+LM Fine-tuning consecutively. A more detailed experimental setup can be found in the appendix. Our model is trained on NVIDIA Tesla A100 (40G). Our code is available at https://anonymous.4open.science/r/lancer-3109/.
@@
We compare the performance of our model to various baselines, including ID-based methods and content-based methods. In addition, we evaluate the performance of recent methods which also use language models as the main model architecture. Specifically, we consider the following baselines.
  • Caser [22] learns local sequential patterns using horizontal and vertical convolutional layers.
  • GRU4Rec [7] uses recurrent neural networks to model interaction history for session-based recommendation.
  • SASRec [9] uses a left-to-right transformer architecture to identify relevant items in historical interactions and in turn predict the next item.
  • BERT4Rec [21] employs bidirectional transformer layers, and captures bidirectional item relationships by training the model to predict masked items at random positions in the sequence.
  • FDSA [30] combines heterogeneous item features using vanilla attention, then separately models item and feature-level sequences using self-attention.
  • MV-RNN [2] employs the RNNs as a sequence encoder to capture user interests from hybrids of item IDs and content features. We implement feature embeddings using the BERT pre-trained model [^5] https://huggingface.co/bert-base-uncased.
  • SASRec-C [9] is a variant of SASRec, in which we directly regard the content feature representation as sequence input. This baseline is an ID-agnostic baseline.

### v4

# Reformulating Sequential Recommendation: Learning Dynamic User Interest with Content-enriched Language Modeling

@@ Abstract
Recommender systems are indispensable in the realm of online applications, and sequential recommendation has enjoyed considerable prevalence due to its capacity to encapsulate the dynamic shifts in user interests. However, previous sequential modeling methods still have limitations in capturing contextual information. The primary reason is the lack of understanding of domain-specific knowledge and item-related textual content. Fortunately, the emergence of powerful language models has unlocked the potential to incorporate extensive world knowledge into recommendation algorithms, enabling them to go beyond simple item attributes and truly understand the world surrounding user preferences. To achieve this, we propose LANCER, which leverages the semantic understanding capabilities of pre-trained language models to generate personalized recommendations. Our approach bridges the gap between language models and recommender systems, resulting in more human-like recommendations. We demonstrate the effectiveness of our approach through a series of experiments conducted on multiple benchmark datasets, showing promising results and providing valuable insights into the influence of our model on sequential recommendation tasks. Furthermore, our experimental codes are publicly available at https://github.com/Gnimixy/lancer.
Keywords:
Sequential Recommendation Language Model.
@@ 1 Introduction
Recommender systems are a crucial engine of various online applications, including e-commerce, advertising, and online videos. They play a pivotal role in discovering user interests and alleviating information overload, thus assisting users in their decision-making process. Over the past years, collaborative filtering (CF) [13, 25, 21] has become one of the most prevalent techniques due to its effectiveness and efficiency in large-scale recommendation scenarios.
@@
By carefully revisiting recent progress in the sequential recommendation, we notice much of the great progress has been driven by achievements in network architecture or pre-training algorithms in natural language processing [14, 22]. For one thing, a large body of advanced architecture models [5, 10] has facilitated the building of high-capacity recommendation models. For another, some researchers attempt to leverage the powerful capacity of language modeling in semantic understanding to perform content-based sequential recommendations [18, 1]. We also notice a few related advancements attempt to integrate pre-trained language models (PLMs) into recommender systems, leveraging their ability to capture semantic information and achieving promising results [8, 11]. However, these works do not depart from the previous paradigm, merely adding a text encoder and continuing to use traditional sequential modeling methods.
@@
With the continuous improvement of model capabilities, we ponder on whether sequential recommendation can be revolutionized to recommend items in a way that mimics human reasoning. Some recent works have attempted to use powerful PLMs as recommender systems by formulating recommendation as a language modeling task, with the item IDs [3] or titles [24] as inputs to the language model. Despite leveraging the inferential capabilities of language models for generating recommendations, these methods yielded unsatisfactory performance, falling behind the previous state-of-the-art results.
@@
In light of the above, language models perform suboptimally to sequential recommendation systems due to their lack of domain-specific knowledge and mastery of item-related textual content. For example, in movie recommendation, The Matrix could be understood by the language model as a mathematical matrix, rather than being associated with hacker-related information, which would significantly reduce the effectiveness of the recommendation. This limitation stems from the absence of side information in sequential recommendation, such as movie genres and synopses, which hinders a comprehensive understanding of user preferences underlying item choices (see Figure 1).
Figure 1: Comparison of sequence recommendation with and without content-enriched information. It can be seen that enriched content information provides more information about the content of the movies, which helps to reduce ambiguity and improve the accuracy of the recommendations.
@@
To address these issues, we propose LANguage modeling for Content-Enriched Recommendation (LANCER) to employ PLMs as sequential recommenders by combining domain-specific knowledge and user behavior. Specifically, we introduce a knowledge prompt that facilitates the acquisition of domain-specific knowledge, and a reasoning prompt that integrates the acquired knowledge and item content information to generate personalized recommendations. To the best of our knowledge, we are the first to integrate enriched content into a language model-based sequential recommender system by transforming the sequential recommendation task into a text generation task.
@@
  • We propose LANCER, which incorporates domain knowledge and item content prompts into PLMs, leveraging their semantic understanding abilities to generate user recommendation results.
  • Our designed model offers the advantage of integrating external knowledge to enhance the modeling of item content understanding. Our approach bridges the gap between language models and recommender systems by fusing abilities from both domains.
@@ 2 Methodology
The proposed LANCER architecture is shown in Figure 2. For each item in the user interaction history, we extract all text content information and input it into a language model with fixed parameters to train the knowledge prompt and obtain a vector representation $h_{i}$ for each item. Afterward, we use an attention mechanism to combine domain knowledge with user interaction histories to form the reasoning prompt, allowing the model to generate the next item that the user may be interested in. Finally, the generated text is mapped to one of the candidate items to be recommended to the user.
2.2 Knowledge Prompt.
When directly applying a language model for item recommendation, a major challenge arises, as language models may not possess the knowledge of the relevant recommendation domains.
@@
To address the above issues, we design continuous prompts to learn recommendation domain information while avoiding catastrophic forgetting that may occur with language model retraining. Inspired by Prefix-tuning [7] and P-tuning [9], we add continuous prefix prompts to guide the PLM without updating the original parameters. We define these as knowledge prompts.
We set a fixed number of knowledge tokens $\varepsilon^{(k)}=[e^{(k)}_{1},...,e^{(k)}_{\theta}]$, where $e^{(k)}_{j}$ represents the $j$-th knowledge token, and $\theta$ denotes the number of knowledge tokens. Using the following formula, we compute the knowledge prompt:
$\Psi^{(k)}=\text{MLP}\left(\text{Emb}(\varepsilon^{(k)})\right),$ (1)
where MLP denotes the multi-layer perceptron, Emb denotes the embedding layer of the language model $\Theta_{f}$ with $L$ layers. During attention computation in the $\ell$-th layer $\ell\in\{1,2,...,L\}$, a distinct prefix of key-value pairs $Z_{\ell,K}$ and $Z_{\ell,V}$ is learnt. $\Psi^{(k)}$ concatenates the parameters of the language model in each layer in the form of $\Psi^{(k)}_{\ell,K}$ and $\Psi^{(k)}_{\ell,V}$. The effect of the knowledge prompt propagates throughout the entire transformer activation layer and subsequently to the following tokens, enhancing them as follows:
$Z_{\ell,K}^{{}^{\prime}}=[\Psi^{(k)}_{\ell,K};Z_{\ell,K}],\quad Z_{\ell,V}^{{}^{\prime}}=[\Psi^{(k)}_{\ell,V};Z_{\ell,V}],$ (2)
@@
where $Z_{\ell,K}^{{}^{\prime}},Z_{\ell,V}^{{}^{\prime}}\in\mathbb{R}^{(\theta+M)\times d}$, $M$ denotes the input length of the language model, and $d$ denotes the dimension of hidden layers. For each item $i$, we utilize the last hidden state $h_{i}$ obtained from the language model’s last input token, considering all preceding tokens. Equation 3 shows how the hidden layer is calculated:
$h_{i}=\Theta_{f}(c_{i}\mid Z_{K}^{{}^{\prime}},Z_{V}^{{}^{\prime}}),$ (3)
@@
where $c_{i}$ denotes the detailed text information of item $i$. Therefore, a user’s behavior sequence can be expressed as $H=[h_{1},...,h_{i},...,h_{n-1}]$. We will pad if the length of the sequence less than the fixed $N$.
At this stage, the training objective focuses on the detailed text information of items. The loss function for $\Theta_{f}$ is depicted in Equation 4, by minimizing the negative log-likelihood of predicting the next token, denoted as $e_{j}$. Training involves backpropagation to update the Emb and MLP corresponding to the knowledge prompts, while the parameters of $\Theta_{f}$ remain frozen.
$\mathcal{L}=\sum_{j}-\log p\left(e_{n}\mid e_{1},...,e_{j-1};\Theta_{LM}\right).$ (4)
2.3 Reasoning Prompt.
With the help of knowledge prompts, we can provide more accurate recommendations for a specific domain. However, this method lacks a modeling of user behavior and a deeper understanding of item information. Yet, user preferences often lie within a more personalized subset of domain information.
@@
To address this issue, we propose reasoning prompts, which employ an attention mechanism to combine domain knowledge and user behavior. In this process, the user’s historical behavioral representation serves as the query, while the domain-related information acts as the key and value. Then we extract the unique reasoning prompt $\Psi^{(r)}$ for each user to help the subsequent model to better provide recommendations, where $A$ is the result calculated by the attention mechanism:
$\Psi^{(r)}=\sigma\left(W_{r}\times A+b_{r}\right).$ (5)
@@
In the generation phase, we use beam search decoding to generate output sequences with the $w$ highest scores, then truncate each output to include only the first new item. Beam search decoding at step $t$ can be expressed as:
$\pi_{t}=\{[x_{t-1},x^{\prime}_{t}]\ |\ x^{\prime}_{t}\in\Omega,x_{t-1}\in\pi_{t-1}\},$ (6)
$\alpha_{t}=\operatorname*{arg\,max}_{\lambda\in\pi_{t},|\lambda|=w}\log P(\lambda\mid\Theta(T,\Psi^{(r)})),$ (7)
where $\Omega$ is the vocabulary set of trainable language model $\Theta$, $w$ is the number of beams, $e^{\prime}_{t}$ denotes a particular text generated at the $t$-th step during beam search decoding, and $\lambda$ is the current set of top-$w$ outputs.
During this phase, the training objective is to enable the language model $\Theta$ to predict text that corresponds to the textual information of the item the user will choose next.
2.4 Item Mapping.
At this point, we have obtained the results of the model prediction, but we still need to map the predicted result to one of all items in the database $\mathcal{D}$. This matching process is conducted as follows: initially, we derive the vector representation sequence of the predicted item $n$, denoted as $\beta=\text{Emb}(\alpha^{(n)})$, where $\beta\in\mathbb{R}^{|\beta|\times d}$ and $|\beta|$ represents the sequence length of $\beta$. Subsequently, we calculate the maximum value for each dimension of the $\beta$ sequence to represent the predicted item as a fixed-length vector, where the maximum value $\gamma_{j}$ for the $j$-th dimension is:
$\gamma_{j}=\max(\beta_{j}),\quad\text{for}\ j=1,2,...,|\beta|.$ (8)
Finally, we compute the cosine similarity (CosSim) in equation 9 between the vector representations of the output $\gamma$ and the titles of each item in the database $\mathcal{D}$, then identify the item with the highest similarity. The representations of all items can be computed only once and used throughout the generation process.
$s_{n}=\mathop{\arg\max}_{\delta\in\mathcal{D}}\textit{CosSim}(\gamma,\delta).$ (9)
@@ 3.1 Experimental Setup
We experimented with three datasets: MovieLens [^1] https://grouplens.org/datasets/movielens/, MIND [^2] https://msnews.github.io/ [19], and Goodreads [^3] https://sites.google.com/eng.ucsd.edu/ucsdbookgraph/home [17]. The detailed statistics of the datasets are presented in Table 1. MovieLens is for movie recommendations, MIND for English news, and Goodreads for books reviews. We organized user data chronologically and limited sequence lengths to 5-10 items, excluding shorter sequences and truncating longer ones.
@@
To evaluate our model, we compute Recall and NDCG for the top 5 and 10 items respectively. The former is an evaluation of un-ranked retrieval sets while the latter reflects the order of ranked lists. We follow the leave one-out strategy described in [5] to split the interactions into training, validation, and testing sets. For the testing stage, we adopt the all-ranking protocol to evaluate recommendation performance.
@@
Dataset Num. User Num. Item Num. Inteaction Sparsity Avg. User Avg. Item
MovieLens 6,040 3,231 72,480 99.63% 12.00 22.43
MIND 91,935 44,908 2,248,027 99.94% 24.45 50.06
Goodreads 120,968 28,480 1,095,758 99.97% 9.06 38.47
@@
We use the GPT-2 [12] pre-trained model from Hugging Face  [^4] https://huggingface.co/gpt2 as our backbone model for experiments. During training, we employ the Adam optimizer with a linear schedule learning rate decay. For different datasets, we keep the following hyperparameters consistent: batch size of 16, maximum token length of 512 for the language model input, maximum input length of 32 tokens for each item (truncated for items exceeding the limit). We employ the Sequential Tuning approach from [6]. Our model is trained on NVIDIA Tesla A100 GPUs with 40GB of memory.
@@
We compare the performance of our model to various baselines, including ID-based and content-based methods. In addition, we evaluate the performance of recent methods which also use PLMs as the main model architecture. Specifically, we compare our model to the following baselines, where Caser [16], GRU4Rec [4], SASRec [5], BERT4Rec [15], S3Rec [26], CL4SRec [20] are ID-based, and , FDSA [23], MV-RNN [2], SASRec+ [5] are content-based.
Table 2: Recommendation performance comparison of different recommendation models over three datasets, in which the best results are in bold and the second best results are underlined. R and N denote Recall and NDCG respectively.
Model MovieLens MIND Goodreads
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
Caser 0.0225 0.0330 0.0151 0.0185 0.0312 0.0575 0.0181 0.0265 0.1058 0.1452 0.0770 0.0898
GRU4Rec 0.0737 0.1316 0.0450 0.0635 0.0981 0.1589 0.0628 0.0816 0.1871 0.2472 0.1394 0.1587
SASRec 0.0967 0.1589 0.0624 0.0823 0.1087 0.1707 0.0714 0.0913 0.2043 0.2641 0.1536 0.1729
BERT4Rec 0.0776 0.1391 0.0466 0.0664 0.0853 0.1359 0.0544 0.0706 0.1892 0.2487 0.1391 0.1583
S3Rec 0.0821 0.1469 0.0509 0.0716 0.0990 0.1586 0.0639 0.0830 0.1753 0.2478 0.1207 0.1442
CL4SRec 0.0996 0.1527 0.0683 0.0853 0.0964 0.1543 0.0625 0.0811 0.1986 0.2599 0.1477 0.1675
FDSA 0.0901 0.1407 0.0595 0.0757 0.1063 0.1728 0.0682 0.0896 0.1713 0.2351 0.1195 0.1400
MV-RNN 0.0853 0.1399 0.0527 0.0703 0.1024 0.1619 0.0656 0.0847 0.1872 0.2486 0.1385 0.1583
SASRec+ 0.0944 0.1454 0.0618 0.0781 0.1107 0.1753 0.0723 0.0930 0.2045 0.2663 0.1533 0.1733
LANCER 0.1086 0.1616 0.0712 0.0884 0.1168 0.1675 0.0766 0.0930 0.2121 0.2673 0.1600 0.1779